the sum of three consecutive odd integers is 89 more than twice the largest integer.find the integer
Found 2 solutions by unlockmath, dabanfield:
Answer by unlockmath(1688)

Hello,
Let's first show x being the first odd integer. x+2 and x+4 the next ones. The equation will look like this:
x+(x+2)+(x+4)=2(x+4)+89 OK tracking so far? Now combine like terms:
3x+6=2x+97 Now subtract 6 from both sides and subtract 2x from both sides to get:
x=91
There we go. 91 in the first odd integer. Plug this into the original to see if it works.
There's a book that might help you out:
www.math-unlock.com

Answer by dabanfield(803)

the sum of three consecutive odd integers is 89 more than twice the largest integer.find the integer
Let x be the smallest of the three integers. The second and third consecutive odd integers would then be x+2 and x+4.
So we have:
x + (x+2) + (x+4) = 2*(x+4)
Solve the above for x and then calculate x+2 and x+4.